# Detectability of Lensed Gravitational Waves in Matched-Filtering Searches

### v1

@@ III The GstLAL search pipeline
In this paper, we will focus on the matched-filtering search employed in the search pipeline, GstLAL. We will focus on the autocorrealtion based signal consistency-test detection statistic, and the ranking statistic assignment of GstLAL, which will aid our discussion in later sections of this paper.
III.1 Autocorrelation Based Signal-Consistency Test
When $\rho(t)$ exceeds a predetermined threshold with $\rho=4$, the search pipeline registers the template parameters as a potential candidate, referred to as a trigger Messick et al. (2017). Non-Gaussianity and “glitches” in the data can also produce high SNR triggers that mimic gravitational waves. GstLAL therefore assigns for each trigger a signal consistency test value $\xi^{2}$, which is based on autocorrelation function. $\xi^{2}$ compares the difference between the expected autocorrelation time series of a template and the actual matched-filtered SNR time series. Mathematically, we have, for a given template $h_{T,\;j}$,
@@ III.1 Autocorrelation Based Signal-Consistency Test
Consider the extreme case where the trigger from a glitch is a scaled delta-function $z_{j}(0)\delta(t)$ at time $t=0$ with amplitude $z_{j}(0)$. The actual SNR time series will then be identically 00 everywhere over the integration range except for $t=0$, which differs from the expected SNR time series (i.e. $z_{j}(0)R_{j}(t)$) significantly. Essentially, the numerator of the $\xi^{2}$ will become large. In fact, the larger the SNR of the noise trigger (i.e. the larger the value of $z_{j}(0)$), the higher the value $\xi^{2}$ will be. Also, a signal that differs from the template waveforms can potentially impact the $\xi^{2}$ and is not trivial Bondarescu et al. (2023).
Figure 2: Time-domain gravitational wave waveforms of a spinless binary black hole with component masses $m_{1}=m_{2}=30M_{\odot}$ in unlensed (black) and lensed (blue) scenarios that represent increasing amplifcation effects. The lensed waveform has the same impact parameter $y=0.01$. Left: redshifted lens mass $M_{\rm{Lz}}=10M_{\odot}$ with match $\mathcal{M}=0.99$. Middle: $M_{\rm{Lz}}=10^{3}M_{\odot}$, $\mathcal{M}=0.97$. Right: $M_{\rm{Lz}}=10^{5}M_{\odot}$, $\mathcal{M}=0.7$. Lensing distortion can enlarge the amplitude.
@@ III.2 Likelihood-Ratio Ranking Statistic
where $\vec{D}_{H}$ are the horizon distances of the detectors, $\vec{O}$ is the set of detectors that registered the trigger, $\vec{\rho}$ are the recorded detector SNRs for the trigger at each participating detector, $\vec{\xi}^{2}$ are the autocorrealtion-based consistency test values computed for each of the participating detectors, $\left[\vec{\Delta t},\vec{\Delta\phi}\right]$ are the sets of differences in times $\Delta t$ and coalescence phases $\Delta\phi$ measured by the participating detectors.
@@
which allows us to compute the False-Alarm-Probability (FAP, or the p-value) as the probability that $M$ independent coincident noise-like events ($\{N_{1},N_{2},...,N_{M}\}$ contains at least one event with a log likelihood ratio $\mathcal{L}$ larger or equal to some threshold $\ln\mathcal{L}^{*}$:
@@
where $e^{-\mathcal{C}(\ln\mathcal{L}^{*}|\text{noise})}$ denotes the probability that a Poisson process with mean rate $\mathcal{C}(\ln\mathcal{L}^{*}|\text{noise})$ can yield an event with log likelihood ratio $\mathcal{L}<\ln\mathcal{L}^{*}$. The FAP denotes the probability that noise can produce a trigger with likelihood ratio $\mathcal{L}$ larger or equal to the likelihood ratio $\mathcal{L}^{*}$ of the trigger under consideration. One would expect triggers from real gravitational-wave signals to have higher likelihood ratios. Hence, the lower the FAP, the more likely the candidate is a gravitational wave. A similar quantity known as the False-Alarm-Rate (FAR), is also evaluated for each candidate. FAR is defined as
@@ IV Brief introduction to Gravitational Lensing
In the following sections, we introduce the theory of gravitational lensing and show how lensing imprints frequency-dependent modulation on the waveforms using the point mass lens model. This modulation includes interference and diffraction, resulting in waveform distortion and amplitude enhancement. We briefly discussed the effects of lensing distortion on detectability in mathced-filtering, using optimal SNR as a proxy.
Figure 3: The role of lensing in the accuracy of the recovered source parameters of lensed gravitational waves. The panels show the histogram of the recovered source parameter with different lens parameters in our first injection campaign (see Tab. 1 and Fig. 2). Left: chirp mass $\mathcal{M}_{c}$. Right: effective spin $\chi_{\rm{eff}}$. The black dotted line represents the injected value of the $30-30$ $M_{\odot}$ spinless binary black hole signals at 125 Mpc. The unlensed injections (blue). Lensed injections with the impact parameter $y=0.01$ and redshifted lens mass $M_{\rm{Lz}}=10M_{\odot}$ with match $\mathcal{M}=0.99$ (red), $10^{3}M_{\odot}$, $\mathcal{M}=0.97$ (green), $10^{5}M_{\odot}$, $\mathcal{M}=0.7$ (purple). Lensing distortion can lead to a substantial reduction in the accuracy of the source parameters.
@@ IV.1 Theory of Gravitational Lensing of Gravitational Waves
Gravitational lensing theory describes how waves propagate in a curved background. When considering the gravitational wave propagation in the weak-field limit, where the Newtonian gravitational potential $U\ll 1$ and and employing the thin-lens approximation, gravitational lensing is well-described by the Fresnel-Kirchhoff diffraction integral over the whole lens plane in frequency domain Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003) :
$F(f,\boldsymbol{y})=\frac{f}{i}\int d^{2}\boldsymbol{x}\leavevmode\nobreak\ e^{i2\pi f\tau_{\rm d}(\boldsymbol{x},\boldsymbol{y})},$ (14)
where $f$ is the frequency of the gravitational waves. $\boldsymbol{x}\equiv(x_{1},x_{2})$ and $\boldsymbol{y}\equiv(y_{1},y_{2})$ represents the image plane and source plane coordinates respectively. The time delay function $\tau_{\rm d}(\boldsymbol{x},\boldsymbol{y})$ is given as
@@
$h_{\rm{L}}(f)\!=\!F(f,y)h_{\rm{UL}}(f)\leavevmode\nobreak\ ,$ (16)
The characteristic wavelength of the CBC sources sensitive the ground-based detector is $\lambda_{\rm{GW}}\sim 10^{6}$ m. Lenses like stars, stellar remnants, star clusters and compact dark halo in a galaxy, have $R_{\rm{Lens}}\sim\lambda_{\rm{GW}}$. This corresponds the redshifted lens mass $M_{\rm Lz}\equiv M_{\rm L}(1+z_{\rm{L}})=[10,10^{5}]M_{\odot}$, where $M_{\rm{L}}$ is the lens mass in the lens frame. In this regime, calculating the amplification factors with Eq. 14 becomes necessary. We refer to it as a wave optics regime in the rest of our text.
@@
$F(w,y)$ $=\exp\bigg{[}\dfrac{\pi w}{4}+i\dfrac{w}{2}\bigg{\{}\text{ln}\bigg{(}\dfrac{w}{2}\bigg{)}-2\phi_{m}(y)\bigg{\}}\bigg{]}$
×$\times\Gamma\bigg{(}1-\dfrac{i}{2}w\bigg{)}{{}_{1}}F_{1}\bigg{(}\dfrac{i}{2}w,1;\dfrac{i}{2}wy^{2}\bigg{)},$ (17)
@@
where $\omega\equiv 8\pi M_{\rm{Lz}}f$ is the dimensionless frequency, , and ${{}_{1}}F_{1}$ is the hypergeometric function  Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003).
@@
We set the ranges of the two lens parameters as $(M_{\rm Lz},y)=([10,10^{5}]M_{\odot},[0.01,1.0])$ to cover regions of long-wavelength regime, wave-dominated regime, geometric-optics regime. For $M_{\rm Lz}<10M_{\odot}$, the resulting lensed waveform is similar to the unlensed waveform. For $M_{\rm Lz}>10^{5}M_{\odot}$, the lensed waveform is in the geometric optics regime, and there are distinct images only with an overall amplification and constant phase shift. Additionally, we exclude cases where $y>1.0$ because lensing distortions are comparatively weaker.
@@
IV.2 The Impact of Lensing on Signal Detectability in Matched-filtering Searches: Optimal SNR
@@
To estimate how lensing within the given parameter space affects the waveform morphology, we calculate the match between the lensed and intrinsic unlensed waveforms. As shown in the left panel of Figure 1, the match in the region with high $M_{\rm Lz}$ and low $y$ values is remarkably low. This occurs because the corresponding lensed signals exhibit distinctive beating patterns due to high magnifications and short time delays, as depicted in Fig. 2. Thus, using unlensed waveforms as templates to analyze the lensed waveforms can lead to inaccurate results.
@@
In matched-filtering searches, the detectability of gravitational-wave signals is approximated by observed SNRs that can be estimated using optimal SNR. In real detection scenarios, detector noise fluctuates over time, which can cause a template with different source parameters, $h_{\rm tem}$, to be the best-fit template for recovering the detected gravitational wave waveform, $h_{\rm true}$. As a result, the observed SNR and optimal SNR values can differ from the values obtained in the absence of noise. In addition to noise effects, lensing (i.e., waveform distortion) can further affect the search pipeline’s ability to recover the best-fit template, resulting in a non-trivial behavior in the observed and optimal SNR. The optimal SNR assumes that the best-fit template is the same as the signal. A higher optimal SNR value generally indicates greater signal detectability. If a signal is lensed by a point mass in a wave optics regime, it is always magnified, resulting in a higher optimal SNR value. The right panel in Fig. 1 illustrates how optimal SNR values increase depending on the lens parameters. As expected, pairs with high $M_{\rm Lz}$ and low $y$, corresponding to higher magnifications, lead to higher optimal SNRs for lensed signals.
@@
Figure 4: The role of lensing in the detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our first injection campaign (see Tab. 1 and Fig. 2) as a function of matched-filter SNR $\rho$ and the signal consistency test $\xi^{2}/\rho^{2}$ (see Eq. 7). The color bar represents the false-alarm-rate (FAR). An event is considered a significant gravitational wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or 1 every 30 days, indicated by the red tick. The circle and triangular markers represent larger and smaller injections than the FAR threshold for the detection, respectively. Upper left: Unlensed injections. The remaining panels are lensed injection with impact parameter $y=0.01$. Upper right: redshifted lens mass $M_{Lz}=10M_{\odot}$ with match $\mathcal{M}=0.99$. Bottom left: $M_{Lz}=10^{3}M_{\odot}$, $\mathcal{M}=0.97$. Bottom right: $M_{Lz}=10^{5}M_{\odot}$, $\mathcal{M}=0.7$. Lensing distortion can introduce a significant drop in $\rho$ and a larger $\xi^{2}$ value in matched-filtering searches

### v2

@@ III The GstLAL search pipeline
In this paper, we will focus on the matched-filtering search employed in the search pipeline, GstLAL. We will focus on the autocorrelation-based signal-consistency test detection statistic, and the ranking statistic assignment of GstLAL, which will aid our discussion in later sections of this paper.
III.1 Autocorrelation-based signal-consistency test
When $\rho(t)$ exceeds a predetermined threshold with $\rho=4$, the search pipeline registers the template parameters as a potential candidate, referred to as a trigger Messick et al. (2017). Non-Gaussianity and “glitches” in the data can also produce high-SNR triggers that mimic gravitational waves. GstLAL therefore assigns for each trigger a signal-consistency test value $\xi^{2}$, which is based on the autocorrelation function. $\xi^{2}$ compares the difference between the expected autocorrelation time series of a template and the actual matched-filtered SNR time series. Mathematically, we have, for a given template $h_{T,\;j}$,
@@ III.1 Autocorrelation-based signal-consistency test
Consider the extreme case where the trigger from a glitch is a scaled $\delta$ function $z_{j}(0)\delta(t)$ at time $t=0$ with amplitude $z_{j}(0)$. The actual SNR time series will then be identically 00 everywhere over the integration range except for $t=0$, which differs from the expected SNR time series [i.e. $z_{j}(0)R_{j}(t)$] significantly. Essentially, the numerator of the $\xi^{2}$ will become large. In fact, the larger the SNR of the noise trigger [i.e. the larger the value of $z_{j}(0)$], the higher the value $\xi^{2}$ will be. Also, a signal that differs from the template waveforms can potentially impact the $\xi^{2}$ and is not trivial Bondarescu et al. (2023).
Figure 2: Time-domain gravitational-wave waveforms of a spinless binary black hole with component masses $m_{1}=m_{2}=30M_{\odot}$ in unlensed (black) and lensed (blue) scenarios that represent increasing amplification effects. The lensed waveform has the same impact parameter $y=0.01$. Left: redshifted lens mass $M_{\rm{Lz}}=10M_{\odot}$ with match $\mathcal{M}=0.99$. Middle: $M_{\rm{Lz}}=10^{3}M_{\odot}$, $\mathcal{M}=0.97$. Right: $M_{\rm{Lz}}=10^{5}M_{\odot}$, $\mathcal{M}=0.7$. Lensing distortion can enlarge the amplitude.
@@ III.2 Likelihood-ratio ranking statistic
where $\vec{D}_{H}$ are the horizon distances of the detectors, $\vec{O}$ is the set of detectors that registered the trigger, $\vec{\rho}$ are the recorded detector SNRs for the trigger at each participating detector, $\vec{\xi}^{2}$ are the autocorrelation-based consistency test values computed for each of the participating detectors, and $\left[\vec{\Delta t},\vec{\Delta\phi}\right]$ are the sets of differences in times $\Delta t$ and coalescence phases $\Delta\phi$ measured by the participating detectors.
@@
which allows us to compute the false-alarm probability (FAP, or the p value) as the probability that $M$ independent coincident noiselike events ($\{N_{1},N_{2},...,N_{M}\}$ contain at least one event with a log likelihood ratio $\mathcal{L}$ larger or equal to some threshold $\ln\mathcal{L}^{*}$:
@@
where $e^{-\mathcal{C}(\ln\mathcal{L}^{*}|\text{noise})}$ denotes the probability that a Poisson process with mean rate $\mathcal{C}(\ln\mathcal{L}^{*}|\text{noise})$ can yield an event with log likelihood ratio $\mathcal{L}<\ln\mathcal{L}^{*}$. The FAP denotes the probability that noise can produce a trigger with likelihood ratio $\mathcal{L}$ larger or equal to the likelihood ratio $\mathcal{L}^{*}$ of the trigger under consideration. One would expect triggers from real gravitational-wave signals to have higher likelihood ratios. Hence, the lower the FAP, the more likely the candidate is a gravitational wave. A similar quantity known as the false-alarm rate (FAR), is also evaluated for each candidate. FAR is defined as
@@ IV Brief introduction to Gravitational Lensing
In the following sections, we introduce the theory of gravitational lensing and show how lensing imprints frequency-dependent modulation on the waveforms using the point mass lens model. This modulation includes interference and diffraction, resulting in waveform distortion and amplitude enhancement. We briefly discussed the effects of lensing distortion on detectability in mathced filtering, using optimal SNR as a proxy.
Figure 3: The role of lensing in the accuracy of the recovered source parameters of lensed gravitational waves. The panels show the histogram of the recovered source parameter with different lens parameters in our first injection campaign (see Table 1 and Fig. 2). Left: chirp mass $\mathcal{M}_{c}$. Right: effective spin $\chi_{\rm{eff}}$. The red solid line on the x-axis represents the injected value of the $30-30$ $M_{\odot}$ spinless binary black hole signals at 125 Mpc. The unlensed injections are blue. Lensed injections with the impact parameter $y=0.01$ and redshifted lens mass $M_{\rm{Lz}}=10M_{\odot}$ with match $\mathcal{M}=0.99$ (red), $10^{3}M_{\odot}$, $\mathcal{M}=0.97$ (green), $10^{5}M_{\odot}$, $\mathcal{M}=0.7$ (purple). Lensing distortion can lead to a substantial reduction in the accuracy of the source parameters.
@@ IV.1 Theory of gravitational lensing of gravitational waves
Gravitational lensing theory describes how waves propagate in a curved background. When considering the gravitational-wave propagation in the weak-field limit, where the Newtonian gravitational potential $U\ll 1$ and employing the thin-lens approximation, gravitational lensing is well described by the Fresnel-Kirchhoff diffraction integral over the whole lens plane in the frequency domain Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003),
$F(f,\boldsymbol{y})=\frac{f}{i}\int d^{2}\boldsymbol{x}\leavevmode\nobreak\ e^{-i2\pi f\tau_{\rm d}(\boldsymbol{x},\boldsymbol{y})},$ (14)
where $f$ is the frequency of the gravitational waves. $\boldsymbol{x}\equiv(x_{1},x_{2})$ and $\boldsymbol{y}\equiv(y_{1},y_{2})$ represent the image plane and source plane coordinates, respectively. The time delay function $\tau_{\rm d}(\boldsymbol{x},\boldsymbol{y})$ is given as
@@
$h_{\rm{L}}(f)\!=\!F(f,\bm{y})h_{\rm{UL}}(f)\leavevmode\nobreak\ ,$ (16)
The characteristic wavelength of the CBC sources sensitive the ground-based detector is $\lambda_{\rm{GW}}\sim 10^{6}$ m. Lenses like stars, stellar remnants, star clusters and compact dark halos in a galaxy, have $R_{\rm{Lens}}\sim\lambda_{\rm{GW}}$. This corresponds to the redshifted lens mass $M_{\rm Lz}\equiv M_{\rm L}(1+z_{\rm{L}})=[10,10^{5}]M_{\odot}$, where $M_{\rm{L}}$ is the lens mass in the lens frame. In this regime, calculating the amplification factors with Eq.( 14) becomes necessary. We refer to it as a wave optics regime in the rest of our text.
@@
$F(w,y)$ $=\bigg{[}\exp\bigg{[}\dfrac{\pi w}{4}+i\dfrac{w}{2}\bigg{\{}\text{ln}\bigg{(}\dfrac{w}{2}\bigg{)}-2\phi_{m}(y)\bigg{\}}\bigg{]}$
×$\times\Gamma\bigg{(}1-\dfrac{i}{2}w\bigg{)}{{}_{1}}F_{1}\bigg{(}\dfrac{i}{2}w,1;\dfrac{i}{2}wy^{2}\bigg{)}\bigg{]}^{*},$ (17)
@@
where $\omega\equiv 8\pi M_{\rm{Lz}}f$ is the dimensionless frequency, and ${{}_{1}}F_{1}$ is the hypergeometric function  Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003).
@@
We set the ranges of the two lens parameters as $(M_{\rm Lz},y)=([10,10^{5}]M_{\odot},[0.01,1.0])$ to cover regions of the long-wavelength regime, wave-dominated regime, and geometric-optics regime. For $M_{\rm Lz}<10M_{\odot}$, the resulting lensed waveform is similar to the unlensed waveform. For $M_{\rm Lz}>10^{5}M_{\odot}$, the lensed waveform is in the geometric-optics regime, and there are distinct images only with an overall amplification and constant phase shift. Additionally, we exclude cases where $y>1.0$ because lensing distortions are comparatively weaker.
@@
IV.2 The impact of lensing on signal detectability in matched-filtering searches: Match and optimal SNR
To estimate how lensing within the given parameter space affects the waveform morphology, we calculate the match between the lensed $\mathcal{M}(h_{\rm{L},}h_{\rm{UL}})$ and intrinsic unlensed waveform and the ratio of the optimal SNR in each lensed and unlensed scenario, $\rho_{\rm{L,opt}}/\rho_{\rm{UL,opt}}$, where $\rho_{\rm{L,opt}}$ is the optimal SNR between the lensed waveforms and $\rho_{\rm{UL,opt}}$ is the optimal SNR between the unlensed waveforms, is considered. A $30$-$30\leavevmode\nobreak\ M_{\odot}$ spinless waveform template is used in both computations. As discussed in Sec. II.2, $\mathcal{M}(h_{\rm{L}},h_{\rm{UL}})$ is normalized by both the lensed and unlensed waveform amplitudes and is hence insensitive to lensing magnification. In contrast, $\rho_{\rm{L,opt}}$ accounts for both waveform distortion and amplitude changes. Hence, if the lensing magnification effect dominates over the waveform distortion effect, we have $\rho_{\rm{L,opt}}/\rho_{\rm{UL,opt}}>1$.
@@ IV.2 The impact of lensing on signal detectability in matched-filtering searches: Match and optimal SNR
As shown in the left panel of Fig. 1, the match in the region with high $M_{\rm Lz}$ and low $y$ values is remarkably low. This occurs because the corresponding lensed signals exhibit distinctive beating patterns due to high magnifications and short time delays, as depicted in the right panel of Fig. 2. Thus, using unlensed waveforms as templates to analyze the lensed waveforms can lead to inaccurate results.
@@
Furthermore, the optimal SNR assumes that the best-fit template is the same as the signal. A higher optimal SNR value generally indicates greater signal detectability. If a signal is lensed by a point mass in a wave optics regime, it is always magnified, resulting in a higher optimal SNR value. The right panel in Fig. 1 illustrates how optimal SNR values increase depending on the lens parameters. As expected, pairs with high $M_{\rm Lz}$ and low $y$, corresponding to higher magnifications, lead to higher optimal SNRs for lensed signals. The right panel of Fig. 2 explains the difference between the left and the right panel in Fig. 1, An isolated point mass can enhance the waveform amplitude by at least five times while $\mathcal{M}(h_{\rm{L},}h_{\rm{UL}})=0.7$. It represents a scenrio where the lensing magnification overwhelms the waveform distortion, result in $\rho_{\rm{L,opt}}/\rho_{\rm{UL,opt}}\sim 10$. This is qualitatively consistent with the matched-filter analysis in Mishra et al. (2024).
To distinguish the lensing distortion from glitches or deviations from general relativity Gupta et al. (2024); Chandramouli et al. (2025); Liu et al. (2024), one needs to perform parameter estimation and Bayesian model analysis, which are beyond the scope of this paper, interested reader please find a systematic study  Gupta et al. (2024); Chandramouli et al. (2025); Liu et al. (2024) to study the systematics between lensing and deviation from general relativity. These analyses are performed based on a confirmed detection from the search pipeline and are therefore subject to bias if the detectability of the signal is not well understood.
In matched-filtering searches, the detectability of gravitational-wave signals is approximated by observed SNRs that can be estimated using optimal SNR. In real detection scenarios, detector noise fluctuates over time, which can cause a template with different source parameters, $h_{\rm tem}$, to be the best-fit template for recovering the detected gravitational-wave waveform, $h_{\rm true}$. As a result, the observed SNR and optimal SNR values can differ from the values obtained in the absence of noise. In addition to noise effects, lensing (i.e., waveform distortion) can further affect the search pipeline’s ability to recover the best-fit template, resulting in a nontrivial behavior in the observed and optimal SNR. Therefore, one need to perform an injection campaign to understand how gravitational lensing influence the detectability in matched-filtering searches.
@@
Figure 4: The role of lensing in the detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our first injection campaign (see Table 1 and Fig. 2) as a function of matched-filter SNR $\rho$ and the signal-consistency test $\xi^{2}/\rho^{2}$ [see Eq.( 7)]. The color bar represents the FAR. An event is considered a significant gravitational-wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or one every 30 days, indicated by the red tick. The circle and triangular markers represent larger and smaller injections than the FAR threshold for the detection, respectively. Upper left: unlensed injections. The remaining panels are lensed injection with impact parameter $y=0.01$. Upper right: redshifted lens mass $M_{Lz}=10M_{\odot}$ with match $\mathcal{M}=0.99$. Bottom left: $M_{Lz}=10^{3}M_{\odot}$, $\mathcal{M}=0.97$. Bottom right: $M_{Lz}=10^{5}M_{\odot}$, $\mathcal{M}=0.7$. Lensing distortion can introduce a significant drop in $\rho$ and a larger $\xi^{2}$ value in matched-filtering searches